Bonnie has been strung up on bungees in the PSL so that her position/orientation can be adjusted however we like. She is now hanging pretty low over the table, rather than being attached to the hanging equipment shelf thing. Butch Cassidy has been hung over the AS table.
Moving Bonnie increased the coherence for the PMC_ERR_F signal, but not the MC_L. Butch Cassidy doesn't have much coherence with either.
I noticed that the coherence would drop off very sharply just after 10 kHz - there would be no further spikes or anything of the sort. I used my computer to play a swept sine wave (sweeping from 20Hz to 10kHz) next to Butch Cassidy to see if the same drop-off occurred in the microphone signal itself. Sure enough, the power spectrum showed a sharp drop around 10kHz. Thinking that the issue was that the voltage dividers had too high impedance, I remade one of them with two 280 Ohm and one 10 Ohm resistor, but that didn't make any difference. So, I'm not sure what's happening exactly. I didn't redo the other voltage divider, so Bonnie is currently not operating.
In her position overlooking whichever table it is that is next to the PSL, Bonnie drummed up some decent coherence with the PSL-PMC_ERR channel, but not so much with the MC_L. I moved her into the PSL itself, and now there is rather good coherence with the PMC_ERR channel, but still not so great for MC_L.
Bonnie's new home in the PSL.
Today I tuned the periscope on the PSL table to align the beam to the Mode Cleaner. With the Wave Front Sensor control off, I minimized the reflection from the MC and maximized the transmission. While doing that I also checked that the transmitted beam after the MC didn't lose the alignment with the interferometer's main Faraday isolator.
In this way, I've got a reflection, as read from the MC_REFLPD_MC, of about 0.6. Then I centered the WFS on the AS table. After that the WFS alignment control brought the reflection to 0.25 and a nice centered bull-eye spot showed on the monitor.
The beam scan (which has been living in the bridge subbasement for a bit now) is in a state of imperfection.
I noticed that:
I took some pictures to try and illuminate the situation - The inverted images are included to make it easier to see the flecks (?) in the slits
I am not sure how to figure out if any bit of the scan is/has been fried.
Pending further investigation, enjoy large error bars in your scan measurements!
PICTURES OF BOTH SLITS ON THE BEAMSCAN HEAD:
After setting up the microphones last week, I modified the Wiener filtering programs so as to include the microphone signals. They didn't seem to do much of anything to reduce the MC_L signal, so I looked at coherences. The microphones don't seem to have much coherence with the MC_L signal at all. I tried moving Bonnie to near the optical table next to the PSL (which isn't in a vacuum, and thus would, presumably, be more affected by acoustic noise), but that didn't seem to make much of a difference. Eventually, I'd like to put a mic in the PSL itself, but I need to work out how to mount it first.
Bonnie's new location.
You can see in bonnie_butch.pdf that none of the mic signals are giving very good coherence, although they all seem to have a peak at 24 Hz. (In fact, everything seems to have a peak there. Must be a resonant frequency of something in the mode cleaner.)
I've also attached plots of the coherences for all six accelerometers and the three Guralp seismometer axes. I plotted the most coherent traces together in the last pdf: the y-axes of the MC2 accelerometer and the two seismometers (the Ranger measures ONLY y) and, interestingly, the z-axis of the MC2 accelerometer. Unsurprisingly, the seismometers are most coherent at the low frequencies, and the MC2-Y accelerometer seems to be coherent at very similar frequencies. The MC2-Z accelerometer, on the other hand, seems to be coherent at the higher frequencies, and is highly complementary to the others. I am not really sure why this would be...
Finally, I was curious about how the noise varies throughout the day, because I didn't want to mistakenly decide that some particular configuration of accelerometers/seismometers/whatever was better than another b/c I picked the wrong time of day to collect the data. So, here is a plot of Wiener filters (using only accelerometer data) taken over 2-hour intervals throughout the entirety of July 6, 2009 (midnight-midnight local).
It's a little bit confusing, and I should probably try to select some representative curves and eliminate the rest to simplify things, but I don't have time to do that before the meeting, so this will have to suffice for now.
Once again, this morning I found the wireless router disconnected from the LAN cable. No martian WiFi was available.
I wonder who is been doing that and for what reason.
The plan for the optical setup has been corrected after it was realized that it would be impossible to isolate a 29.501 MHz frequency from a 29.499 MHz one because they are so close in value. Instead, we decided to adopt the setup pictured below. In this way, the low-pass filter should have no trouble isolating 29.501-29.5 MHz from 29.501 + 29.5 MHz. Also, we decided to scrap the idea of sending Alberto's laser through a fiber optic cable after hearing rumors of extra lasers. Since I shouldn't have to share a beam when the second laser comes in, I plan on setting up both lasers on the same optics bench. I've been working on the software while waiting for supplies, but I should be able to start building the trigger box today (assuming the four-pair cable is delivered).
The last week I've spent mostly working on calculating shot noise and other sensitivities in three michelson sensor setups, the standard michelson, the "long range" michelson (with wave plates), and the proposed EUCLID setup. The goal is to show that there is some inherent advantage to the latter two setups as displacement sensors. This involved looking into polarization and optics a lot more, so I've been spending a lot of time on that also. For example, the displacement sensitivity/shot noise on the standard michelson is around 6:805*10^-17 m/rHz at L_=1*10^-7m, as shown in the graph.
This week, I've been working on adapting the last week's circuit to make it buildable. Mostly this has involved picking components that are already in the lab, adding tunable components when necessary, and planning roughly how the components should be laid out on a board. I then built the circuit and put it in a box with BNC connectors for easy connection during testing. A picture of the built circuit is attached.
For initial testing, the transformer was removed from the design; since this changed the response of the circuit, I added two resistors to correct the response. A figure showing a schematic of the built circuit is attached. The expected responce of the circuit is also shown; the magnitude (solid) and phase (dashed) of the voltage across the EOM are shown in green, and the impedance of the circuit is shown in blue. While this response has sharp peaks and 50 Ohms (34 dB) of impedance at resonances, the gain is low compared to the circuit with the transformer. This means that, as is, this circuit cannot be used to drive the EOM; it is simply for testing purposes.
I am attempting to use the DTT program to look at the coherence of the individual accelerometer signals with the MC_L signal. Rana suggested that I might break up the XYZ configuration, so i wanted to see how the coherence changed when I moved things around over the past couple of weeks, but I keep getting a synchronization error every time I try to set the start time to more than about 3 days ago. I tried restarting the program and checking the "reconnect" option in the "Input" tab, neither of which made any kind of difference. I can access this data with no problem from the Data Viewer and the Matlab scripts, so I'm not really sure what is happening. Help?
EDIT: Problem solved - Full data was not stored for the time I needed to access it for DTT.
Alan and Alberto conducted a tour of 40 high-school students.
It may be the same tour that Rana found a spare PMC during the tour explanation as far as I remember...
Joe, Alberto and Steve
We tested gate valve V1 interlock by :
1, decelerated rotation by brake from maglev controller unit.
2, turned maglev controller off from controller unit.
3, unpluged 220VAC plug from wall socket
None of the above action triggered V1 to close. This needs to be corrected in the future.
The MEDM monitor screen of maglev indicated the correct condition changes.
I hooked up Bonnie and Clyde last night and tested it today. First I tried some loud noises to make sure I could identify them on the readout. Then, Steve suggested I try to look for some periodic stuff. I set up Butch Cassidy and the Sundance Kid on the cabinets by the MC2 optic. Now for graphs!
I tapped on the microphone a few times. I also yelled a bit, but this is sampling by seconds, so perhaps they got overwhelmed by the tapping.
This time I tried some more isolated yells. I started with a tap so I'd be sure to be able to recognize what happened. Apparently, not so necessary.
Here, it looks like a pretty strong periodic pattern on the second mic (Butch Cassidy). I replaced the lines with dashed ones where the pattern was a little less clear. Possibility interference from something. Mic1 (Bonnie) seems to show a pretty regular beat pattern, which seems reasonable, as it isn't particularly close to any one instrument fan.
So, anyway. I thought those were neat. And that I wanted to share.
I clamped Bonnie (microphone) to the top of a chamber near the vertex of the arms and placed Clyde (pre-amp) on the table right below (see picture). The cable was laid and Bonnie and Clyde are plugged into port #13 on the ADC. The second cable was plugged into port #14, but it is not connected to anything. I placed the looped up cable on top of the cabinet holding the ADC.
Note: the angle in the photograph is such that we are looking along the y-arm.
I tested the voltage dividers and was getting up to about 3V. I retested the mic w/o the voltage divider in place, and, lo and behold, I was able to generate about 70-75V (previously, I maxed out at 45V). I'm not 100% sure why this was, but it occurs to me that, before, the sounds I was generating were short in duration (loud claps, short yelps). This time, I tried yelling continuously into the microphone. So, probably, I simply wasn't seeing the real peak before on the scope because it was too short to pick up. I have corrected the voltage dividers (by replacing the 25.5 kOhm resistors, which were in parallel with the ADC, with 10 kOhm resistors, taking the voltage ratio to ~60:1) and tested them. I haven't been able to generate more than 1500 mV, so I think they are safe. (It's possible we would have been fine with the old setup, since I think it would be hard to get any noises as loud as I was making, but better safe than sorry, right?)
I'm attaching a diagram of the new-and-improved voltage dividers.
This past week, I have building a sine wave rectifier and trying to write a simple program that displays a ccd image to familiarize myself with the code. I also wrote a progress report in which I included the following images of the sine wave rectifier circuit as well as the optical chain including the phase-locked loop. The hirose connector arrived so I can begin soldering the electronics together and testing the trigger box with the ccd. I am waiting on the universal PDH box as well as another fiber coupler to begin setting up the optics. In order to avoid the frustrations associated with sending a laser beam down a long pipe to an optical bench across the room, I will be transmitting laser 1 to the ccd by means of a fiber optic cable and dealing with the alternative new and exciting frustrations.
Since last week, I have come up with a new circuit, which is shown in the attached figure. The magnitude (solid) and phase (dashed) of the voltage across the EOM (red), the ratio between the voltage across the EOM and the voltage across the primary nodes of the transformer (blue), and the impedance through the primary port of the transformer (red) are also shown in an attached figure. As can be seen on the plot, resonance occurs at 11 MHz, 29.5 MHz, and 55 MHz, as specified. Also, at these resonant frequencies, the impedance is about 50 Ohms (34 dB). The gain between the voltage across the EOM and the voltage across the primary nodes of the transformer (or output of the crystal oscillator) is about 12 dB; we'd like a higher gain than this, but this gain is primarily governed by the ratio between the secondary and primary inductances in the transformer, and we are using the largest available ratio (on the Coilcraft website) that has the necessary bandwidth. Because of this, we will likely have to add another component between the crystal oscillator and the EOM circuit, to get the voltage to the desired 8.5 Vp across the EOM (for an optical modulation depth of 0.1 rad).
The current and power through the primary port of the tranformer are 43-85 mA and 25-92 mW, respectively. Since the transformer ratings are 250 mA and 1/4 W for current and power; these values should be safe to use with the intended transformer. Also, the highest power dissipated by a resistor in the circuit (not including the 50 Ohm resistor that is part of the crystal oscillator setup) is around 74 mW.
I spent the last week working a lot with the differences between a basic Michelson readout and the new one as a displacement sensor. The new one (w/ wave plates) ends with two differently polarized beams and should have better sensitivity; I've also been going through noise/sensitivity calculations for each, although that hit a road block when I had to start the 1st SURF progress report, which has taken up most of my time since Saturday.
Tonight I tried to lock the interferometer. At the first attempts the arm power didn't go above about 4. The mode cleaner seemed to be not well aligned and it lost lock or got stuck on a wrong mode. I had to run the MC_UP and MC_DOWN scripts to lock it again.
After that the locking proceed more smoothly; at least till a power level in the arms of about 60. Then again the mode cleaner lost lock and I had to run the scripts again. Without the MCWFS servo off the MC reflected power is still rather high (about 1.7); also even when the WFS servo is engaged the reflected power is about 0.5, versus 0.3 that it should be.
Those are both signs of a not very good alignment. Tomorrow I'll have to work on the injection periscope on the PSL table to try to fix that.
So, I'm double-posting, but I figured the last post was long enough as it was, and this is about something different. After double and triple checking the XLR cables, I hooked up the microphone setup (mic---preamp---output) to the oscilloscope to figure out what kind of voltage would register with loud noises. So, I clapped and shouted and forgot to warn the other people in the lab what I was doing (sorry guys) and discovered that, even on the lowest gain setting, my loud noises were generation 2-3 times as much voltage as the ADC can handle (2V). And, since our XLR cables are so freaking long, we probably want to go for a higher gain, which puts us at something like 20 times too much voltage. I doubt this is really necessary, but it's late (early) and I got camera-happy, so I'm going to share anyway:
So, to deal with this issue, I made some nifty voltage dividers. Hopefully they are small enough to fit side-by-side in the ports without needing extra cableage. Anyway, they should prevent the voltage from getting larger than 2V at the output even if the mic setup is producing 50V. Seeing as my screaming as loud as I could about 2mm away from the mic at full gain could only produce 45V, I think this should be pretty safe. I put the ADC in parallel with a 25.5 kOhm resistor, which should have a noise like 10^-8 V/rHz. This is a lot smaller than 1 uV/rHz (the noise in the ADC, if I understood Rana's explanation correctly), so the voltage dividers should pose a noise issue. Now for pictures.
I opened one so you can see its innards.
In case the diagram on the box was too small to decipher...
And finally, I came up with a name scheme for the mics and pre-amps. We now have two Bluebird (bacteriophage) mics named Bonnie and Butch Cassidy. Their preamps are, naturally, Clyde and The Sundance Kid. Sadly, no photos. I know it's disappointing. Also, before anyone gives me crap for putting the labels on the mics upside-down, they are meant to be hung or mounted from high things, and the location (and stiffness) of the cable prevents us from simply standing them up. So they will more than likely be in some kind of upsidedownish position.
I have been working on finding the best spots to put the accelerometer sets in order to best subtract out noise (seismometers next!). Here is a plot of what I've done so far:
All of these were 80-minute samples. The dashed line is unfiltered, solid line filtered. So, Setup #1 looks the best so far, but I didn't leave it there very long, so perhaps it was just a really awesome 80 minutes. I've put the accelerometers back in the Setup #1 position to make sure that it is really better.
And, in case you can't intuitively figure out what configuration the accelerometers are in by such descriptive names, here are some helpful pictures. I didn't know about the digital cameras at first, so these are actually sketches from my notebook, which I helpfully labeled with the setup numbers, color-coded to match the graph above! Also, there are some real-life photographs of the current arrangement (Setup #1' if you forgot).
Doesn't this one look kind of Quentin Blake-esque? (He illustrated for Roald Dahl.)
This is the MC1 set.
Guess which one this is!
[ Jenne, Clara ]
We made new channels for the microphones which came in this week, by editing C1ADCU_PEM.ini (and making an appropriate backup before we modified it) then restarting the framebuilder and the frontend computer C0DCU1. The new channels are:
These are connected to channels 13 and 14 on the PEM ADCU board, just next to the GURALP seismometer channels.
Clara is testing the mics so the max output voltage can be limited to +-2V for the DAQ, then we'll hook them up to our new channels and listen to the IFO (and all the audio frequency noises around it).
We now have two 80-foot, female-to-female XLR cables for our pretty new microphones, one yellow and one purple. They have been tested and appropriately labeled.
Also, here is a very helpful pdf for how to properly attach the XLR connectors to a raw quad cable, as well as one for how to put the actual connectors together (ignore the cable instructions on the connector page... the cable depicted is not a quad cable).
I moved the MC1 set of accelerators. Might have bumped things. If things aren't working, look around the MC1 chamber.
Also, I constructed two new XLR cables, but have not tested them yet.
The 40m lab specific safety training is done. The participants were
Stephanie Erickson, Clara Bennett, Chris Zimmerman, Zach Commings, Michelle Stephen surfs and Drew Cappel postock.
They have already went through the Caltech Safety Office laser and general safety training.
They still have to read, understand and sign the the SOP for the laser & lab
tdsresp is broken on our linux control room machines. I made a little perl replacement which uses the DiagTools.pm perl module, called pzresp. It's in the $SCRIPTS/general directory, and so in the path of all the machines. I also edited the cshrc.40m file so that on linux machines tdsresp points to this perl replacement.
I've patched DiagTools.pm to circumnavigate the tdsdmd bug described here. I also added a function to DiagTools.pm called diagRespNoLog, which is just like diagResp but without that pesky log file.
Here's the output from the tdsresp binary on CentOS:
allegra:~>tdsresp 941.54 10000 100 10 C1:LSC-ITMX_EXC C1:LSC-PD1_Q C1:LSC-PD1_I
nan nan nan nan nan nan nan
nan nan nan nan nan nan nan
nan nan nan nan nan nan nan
*** glibc detected *** tdsresp: free(): invalid next size (fast): 0x089483e8 ***
======= Backtrace: =========
======= Memory map: ========
00242000-00249000 r-xp 00000000 fd:00 15400987 /lib/librt-2.5.so
00249000-0024a000 r--p 00006000 fd:00 15400987 /lib/librt-2.5.so
0024a000-0024b000 rw-p 00007000 fd:00 15400987 /lib/librt-2.5.so
009f9000-00a13000 r-xp 00000000 fd:00 15400963 /lib/ld-2.5.so
00a13000-00a14000 r--p 00019000 fd:00 15400963 /lib/ld-2.5.so
00a14000-00a15000 rw-p 0001a000 fd:00 15400963 /lib/ld-2.5.so
00a17000-00b55000 r-xp 00000000 fd:00 15400974 /lib/libc-2.5.so
00b55000-00b57000 r--p 0013e000 fd:00 15400974 /lib/libc-2.5.so
00b57000-00b58000 rw-p 00140000 fd:00 15400974 /lib/libc-2.5.so
00b58000-00b5b000 rw-p 00b58000 00:00 0
00b5d000-00b70000 r-xp 00000000 fd:00 15400984 /lib/libpthread-2.5.so
00b70000-00b71000 r--p 00012000 fd:00 15400984 /lib/libpthread-2.5.so
00b71000-00b72000 rw-p 00013000 fd:00 15400984 /lib/libpthread-2.5.so
00b72000-00b74000 rw-p 00b72000 00:00 0
00b76000-00b78000 r-xp 00000000 fd:00 15400981 /lib/libdl-2.5.so
00b78000-00b79000 r--p 00001000 fd:00 15400981 /lib/libdl-2.5.so
00b79000-00b7a000 rw-p 00002000 fd:00 15400981 /lib/libdl-2.5.so
00b7c000-00ba1000 r-xp 00000000 fd:00 15400975 /lib/libm-2.5.so
00ba1000-00ba2000 r--p 00024000 fd:00 15400975 /lib/libm-2.5.so
00ba2000-00ba3000 rw-p 00025000 fd:00 15400975 /lib/libm-2.5.so
00bca000-00bdd000 r-xp 00000000 fd:00 15401011 /lib/libnsl-2.5.so
00bdd000-00bde000 r--p 00012000 fd:00 15401011 /lib/libnsl-2.5.so
00bde000-00bdf000 rw-p 00013000 fd:00 15401011 /lib/libnsl-2.5.so
00bdf000-00be1000 rw-p 00bdf000 00:00 0
00dca000-00dd5000 r-xp 00000000 fd:00 15400986 /lib/libgcc_s-4.1.2-20080825.so.1
00dd5000-00dd6000 rw-p 0000a000 fd:00 15400986 /lib/libgcc_s-4.1.2-20080825.so.1
08048000-080b7000 r-xp 00000000 00:17 6455328 /cvs/cds/caltech/apps/linux/tds/bin/tdsresp
080b7000-080ba000 rw-p 0006e000 00:17 6455328 /cvs/cds/caltech/apps/linux/tds/bin/tdsresp
080ba000-080bb000 rw-p 080ba000 00:00 0
0893d000-0896b000 rw-p 0893d000 00:00 0 [heap]
f5e73000-f5e74000 ---p f5e73000 00:00 0
f5e74000-f6874000 rw-p f5e74000 00:00 0
f692d000-f6931000 r-xp 00000000 fd:00 15400995 /lib/libnss_dns-2.5.so
f6931000-f6932000 r--p 00003000 fd:00 15400995 /lib/libnss_dns-2.5.so
f6932000-f6933000 rw-p 00004000 fd:00 15400995 /lib/libnss_dns-2.5.so
f6956000-f6a12000 rw-p f6a31000 00:00 0
f6a74000-f6a7d000 r-xp 00000000 fd:00 15400997 /lib/libnss_files-2.5.so
f6a7d000-f6a7e000 r--p 00008000 fd:00 15400997 /lib/libnss_files-2.5.so
f6a7e000-f6a7f000 rw-p 00009000 fd:00 15400997 /lib/libnss_files-2.5.so
f6a7f000-f6a80000 ---p f6a7f000 00:00 0
f6a80000-f7480000 rw-p f6a80000 00:00 0
f7480000-f7481000 ---p f7480000 00:00 0
f7481000-f7e83000 rw-p f7481000 00:00 0
f7e83000-f7f63000 r-xp 00000000 fd:00 6236924 /usr/lib/libstdc++.so.6.0.8
f7f63000-f7f67000 r--p 000df000 fd:00 6236924 /usr/libAbort
I spent the week reading up on filter algorithm theory, particularly Wiener filtering. I have also learned how to get data from specific channels at specific times, and I've been getting myself acquainted with Matlab (which I have not previously used). Finally, I started messing around with the positioning of the accelerometers and seismometers in order to try to find the setup that yields the best filtration.
This week, I've been reading some literature concerning PLL and familiarizing myself with LINUX, MATLAB, and high-pass filter circuits. In MATLAB, I started constructing matrices to be used for a beam path analysis from the laser output to the ccd camera. I also built a simple high-pass filter on a bread-board that Joe and I are currently testing with the spectrum analyzer.
When I said "MC1/MC2 accelerometers," I meant the entire three-axis accelerometer set at each point.
I have created the attached EOM circuit with resonances at 11 MHz, 29.5 MHz, and 55 MHz (the magnitude and phase of the voltage across the EOM are shown in the attached plot). The gain is roughly the same for each resonant peak. Although I have managed to get the impedances at all of the resonant frequencies to equal each other, I am having more trouble getting the impedances to be 50 Ohms (they are currently all around 0.66 Ohms).
For the current circuit, initial calculations show that we will need around 4.7 - 14.2 A of current to drive the EOM at the desired voltage (8 - 24 V); this is much higher than the current rating of most of the available transformers (250 mA), but the necessary current will change as the impedance of the circuit is corrected, so this is probably not a cause for concern. For example, the necessary driving voltages for the current circuit are (2.8 - 8.5 V); if we assume that the 50-Ohm impedance will be purely resistive, then we get a current range of 56 - 170 mA.
I've spent most of the last week doing background reading; fourier transforms, shm, e&m, and other physics that I didn't cover at school. I also read a few chapters in Saulson, especially the chapter on noise and shot noise. To get a better grip on what I'm going to be doing I read through the polarization chapter in Hobbs' "Optics" text, mostly on wave plates since that's a large part of this readout. Since then I've been working up to calculating the shot noise, starting with the electric field throughout the new interferometer readout.
The Maglev is running for 10 days with V1 closed. The pressure at the RGA-region is at 2e-9 torr on CC4 cold cathode gauge.
Valve VM2 to Rga-only was opened 6 days ago. The foreline pressure is still 2.2e-6 torr with small Varian turbo ~10 l/s on cc2
Daily scans show small improvement in large amu 32 Oxygen and large amu 16, 17 and 18 H20 water peaks.
Argon calibration valve is leaking on our Ar cylinder and it is constant.
The good news is that there are no fragmented hydrocarbons in the spectrum.
The Maglev is soaked with water. It was seating in the 40m for 4 years with viton o-ring seals
However I can not explan the large oxygen peak, either Rai Weiss can not.
The Maglev scans are indicating cleanliness and water. I'm ready to open V1 to the IFO
V1 valve is open to IFO now. V1 interlock will be tested tomorrow.
Valve configuration: VAC NORMAL with CRYO and Maglev are both pumping on the IFO
The IFO RGA scan is normal.
The Cryo needs to be regenerated next. It has been pumping for 36 days since last regenerated.
This has to be done periodically, so the Cryo's 14 K cold head is not insulated by by ice of all things pumped away from the IFO
Both accelerometers have been moved in an attempt to optimize their positions. The MC1 accelerometer was moved from one green bar to the other (I don't know what to call them) at the base of the MC1 and MC3 chambers. That area is pretty tight, as there is an optical table right there, and I did my best to be careful, but if you suspect something has been knocked loose, you might check in that area. The MC2 accelerometer was moved from the horizontal bar down to the metal table on which the MC2 chamber rests.
Sun Jun 21 00:06:43 PDT 2009
Mar 6 15:46:32 nodus sshd: [ID 800047 auth.crit] fatal: Timeout before authentication for 220.127.116.11
Mar 10 11:11:32 nodus sshd: [ID 800047 auth.crit] fatal: Timeout before authentication for 18.104.22.168
Mar 11 13:27:37 nodus sshd: [ID 800047 auth.error] error: connect_to 22.214.171.124 port 7000: Connection refused
Mar 11 13:27:37 nodus sshd: [ID 800047 auth.error] error: connect_to nodus port 7000: failed.
Mar 11 13:31:40 nodus sshd: [ID 800047 auth.error] error: connect_to 126.96.36.199 port 7000: Connection refused
Mar 11 13:31:40 nodus sshd: [ID 800047 auth.error] error: connect_to nodus port 7000: failed.
Mar 11 13:31:45 nodus sshd: [ID 800047 auth.error] error: connect_to 188.8.131.52 port 7000: Connection refused
Mar 11 13:31:45 nodus sshd: [ID 800047 auth.error] error: connect_to nodus port 7000: failed.
Mar 11 13:34:58 nodus sshd: [ID 800047 auth.error] error: connect_to 184.108.40.206 port 7000: Connection refused
Mar 11 13:34:58 nodus sshd: [ID 800047 auth.error] error: connect_to nodus port 7000: failed.
Mar 12 16:09:23 nodus sshd: [ID 800047 auth.crit] fatal: Timeout before authentication for 220.127.116.11
Mar 14 20:14:42 nodus sshd: [ID 800047 auth.crit] fatal: Timeout before authentication for 18.104.22.168
Mar 25 19:47:19 nodus sudo: [ID 702911 local2.alert] controls : 3 incorrect password attempts ; TTY=pts/2 ; PWD=/cvs/cds ; USER=root ; COMMAND=/usr/bin/rm -rf kamioka/
Mar 25 19:48:46 nodus su: [ID 810491 auth.crit] 'su root' failed for controls on /dev/pts/2
Mar 25 19:49:17 nodus last message repeated 2 times
Mar 25 19:51:14 nodus sudo: [ID 702911 local2.alert] controls : 1 incorrect password attempt ; TTY=pts/2 ; PWD=/cvs/cds ; USER=root ; COMMAND=/usr/bin/rm -rf kamioka/
Mar 25 19:51:22 nodus su: [ID 810491 auth.crit] 'su root' failed for controls on /dev/pts/2
Jun 8 16:12:17 nodus su: [ID 810491 auth.crit] 'su root' failed for controls on /dev/pts/4
12:06am up 150 day(s), 11:52, 1 user, load average: 0.05, 0.07, 0.07
12:06am up 150 day(s), 11:52, 1 user, load average: 0.05, 0.07, 0.07
Today I found the elog down, so I rebooted it following the instructions in the wiki.
I have the impression that Nodus has been rebooted since last night, hasn't it?
I measured the scatter from the eLIGO beam dumps as best I could. The experiment setup is shown in the attached diagram.
After familiarizing myself with the equipment in the morning I noticed three issues with the setup
1 - around the minimum scatter the back scatter from the beam dump is very susceptible to the incident angle (makes sense since the Si plate inside the beam dump at Brewster's angle when there is minimum scatter).
2 - The mirrored plug (Part 20 in D0900095) which is suppose to be used for alignment is not very effective. It moves around too much in its hole in the front face of the beam dump. Just by touching it I could make the reflected beam jump around by about 0.1 radians.
- I think to align these properly we'll have to partly assemble the dumps. If we leave off the front plate of the horn then we can measure the reflection off the Si. If we measure this with a power meter then alignment becomes a simple matter of rotating until this reflection is minimized.
3. - For this measurement the incident beam was a small (~ 1mm diameter) central beam with a small amount of spray of laser light beyond that central region. This spray was hitting the aluminium front face of the beam dump and was scattering back to the photodiode. This was clearly the limiting factor in the measurement. Most of this light was spread horizontally so I placed a couple of pieces of black glass on either side of the aperture, just blocking the edges a little. This reduce the background reading at the minimum scatter from 17.0uV to around 4.5uV with still a little bit of light hitting the top and bottom of beam dump face.
The incident power on the beam dump fluctuated a little but was in the range 20.5 to 22mW. The response of the PD is approximately 0.2 A/W and the transimpedance is 7.5E4 V/A.
The SR830 Sensitivity was set to 1x1 mV.
It was difficult to measure the actual angle of incidence. The dump pivoted about a point directly under the input aperture at the front. By measuring the displacement of a point on the back of the dump as I rotated it and knowing the distance between this point and the pivot point I was able to make a reasonably accurate measurement of a range of angles about the minimum.
The measured scatter (in V measured directly by the PD and as a fraction of the incident power) is shown in the attached plots.
I think I can do a better job cleaning up the incident beam - so these numbers only represent an upper limit on the scatter.
attachment 1: beam dump assembly
attachment 2: experimental layout
attachment 3: scatter measurement
attachment 4: BRDF - (scatter divided by the solid angle = 1.1 m steradians)
attachment 5: (slightly blurred )photo of dump - overhead view
This is a ratio of PD1_I to PD1_Q (so a ratio of the two quadratures of AS166), measured in an anti-spring state. It's not flat because our set up has single sideband RF heterodyne detection, and using a single RF sideband as a local oscillator allows one to detect different quadratures by using different RF demodulation phases. So the variation in frequency is actually a measure of how the frequency response of DARM changes when you vary the detection quadrature. This measure is imperfect because it doesn't account for the effect of the DARM loop.
Even though you can choose your detection quadrature with this setup, you can't get squeezed quantum noise with a single RF sideband. The quantum noise around the other (zero-amplitude) RF sideband still gets mixed in, and negates any squeezing benefits.
Here's a noise spectrum of the RSE interferometer, in anti-spring mode, with RF readout. I'd say the calibration is "loose."
I used the Buonanno & Chen modification of the KLMTV IFO transfer functions to model the DARM opto-mechanical response. I just guessed at the quadrature, and normalized the optical gain at the frequency of the calibration line used (927Hz, not visible on the plot).
I checked out a copy of matapps into /cvs/cds/caltech/apps/lscsoft so that I could find the matlab function strassign.m, which is necessary for some old mDV commands to run. I don't know why it became necessary or why it disappeared if it did.
For the 40m Upgrade, we plan to eliminate the Mach-Zehnder and replace it with a single EOM driven by all three modulation frequencies that we'll need: f1=11MHz, f2=5*f1=55MHz, fmc=29.5MHz.
A frequency generator will produce the three frequencies and with some other electronics we'll properly combine and feed them to the EOM.
The frequency generator will have two crystals to produce the f1 and fmc signals. The f2 modulation will be obtained by a frequency multiplier (5x) from the f1.
The frequency multiplier, for the way it works, will inevitably introduce some unwanted harmonics into the signals. These will show up as extra modulation frequencies in the EOM.
In order to quantify the effects of such unwanted harmonics on the interferometer and thus to let us set some limits on their amplitude, I ran some simulations with Optickle. The way the EOM is represented is by three RF modulators in series. In order to introduce the unwanted harmonics, I just added an RF modulator in series for each of them. I also made sure not to leave any space in between the modulators, so not to introduce phase shifts.
To check the effect at DC I looked at the sensing matrix and at the error signals. I considered the 3f error signals that we plan to use for the short DOFs and looked at how they depend on the CARM offset. I repeated the simulations for several possible amplitude of the unwanted harmonics. Some results are shown in the plots attached to this entry. 'ga' is the amplitude ratio of the unwanted harmonics relative to the amplitude of the 11 & 55 MHz modulations.
Comparing to the case where there are no unwanted harmonics (ga = 0), one can see that not considerable effect on the error signals for amplitudes 40dB smaller than that of the main sidebands. Above that value, the REFL31I signals, that we're going to use to control PRCL, will start to be distorted: gain and linearity range change.
So 40 dB of attenuation in the unwanted harmonics is probably the minimum requirement on the frequency multiplier, although 60dB would provide a safer margin.
I'm still thinking how to evaluate any AC effect on the IFO.
** TODO: Plot DC sweeps with a wider range (+/- 20 pm). Also plot swept sines to look for changes in TFs out to ~10 kHz.
With the common mode servo bandwidth above 30kHz and the BOOST on (1), I was able to switch on the test mass dewhitening. Finally.
Last night Rob ran senseDRM and loadDRMImatrixData and came up with the following for the input matrix:
tdswrite C1:LSC-ITMTRX_b2 0.065778 \
C1:LSC-ITMTRX_d2 2.2709 \
C1:LSC-ITMTRX_f2 2.9361 \
C1:LSC-ITMTRX_122 0.42826 \
C1:LSC-ITMTRX_b3 -0.064839 \
C1:LSC-ITMTRX_d3 -0.016913 \
C1:LSC-ITMTRX_f3 -0.021576 \
C1:LSC-ITMTRX_123 -0.0025243 \
C1:LSC-ITMTRX_b5 0.3719 \
C1:LSC-ITMTRX_d5 1.3109 \
C1:LSC-ITMTRX_f5 -0.16412 \
C1:LSC-ITMTRX_125 0.39574 \
C1:LSC-ITMTRX_33 0 \
C1:LSC-ITMTRX_42 0 \
Today, I reran these and got the following, and DD_handoff remained happy:
tdswrite C1:LSC-ITMTRX_b2 -0.10329 \
C1:LSC-ITMTRX_d2 2.0344 \
C1:LSC-ITMTRX_f2 3.2804 \
C1:LSC-ITMTRX_122 0.22516 \
C1:LSC-ITMTRX_b3 -0.076292 \
C1:LSC-ITMTRX_d3 -0.014603 \
C1:LSC-ITMTRX_f3 -0.12101 \
C1:LSC-ITMTRX_123 0.0054128 \
C1:LSC-ITMTRX_b5 0.33521 \
C1:LSC-ITMTRX_d5 1.1425 \
C1:LSC-ITMTRX_f5 -0.32759 \
C1:LSC-ITMTRX_125 0.25877 \
C1:LSC-ITMTRX_33 0 \
C1:LSC-ITMTRX_42 0 \
I wanted to remeasure with the canonical output matrix (-0.7 from MICH to PRM and 0.7 from MICH to SRM), but the DRM freaked out when MICH to PRM went below -0.3.